# CMHSU: An R Statistical Software Package to Detect Mental Health Status, Substance Use Status, and their Concurrent Status in the North American Healthcare Administrative Databases

### v1

@@ 1.2 North American Healthcare Administrative Databases
Country Level Hospital Data Physician Data Canada Federal Discharge Abstract Database (DAD) National Ambulatory Care Reporting System (NACRS) CIHI Physician Claims Data Canadian MIS Database (CMDB) Province MED-ÉCHO (Quebec) data Ontario Health data Alberta Health data Medical Services Plan (MSP) Data (BC) Alberta Health Physician Claims Régie de l’assurance maladie du Québec (RAMQ) Ontario Health Insurance Plan (OHIP) Territory Yukon Hospital Corporation Dis-charge Data Northwest Territories Health Authority Records Yukon Health Care Insurance Plan Claims Northwest Territories Medical Billing Data United States Federal National Inpatient Sample (NIS) Medicare Provider Analysis and Review (MEDPAR) Veterans Affairs (VA) Inpatient Data Medicare Physician/Supplier Part B Claims Data National Ambulatory Medical Care Survey (NAMCS) Veterans Affairs (VA) Outpatient Data State State Inpatient Databases (SID) California OSHPD New York SPARCS State Medicaid Claims Data All-Payer Claims Databases (APCDs) Territory Guam Memorial Hospital Authority Records American Samoa Dept. of Health Inpatient Data Guam Medicaid and CHIP Claims Data American Samoa Medicaid Claims1
@@ 2.1 The Background and Installation
(accessed on 10 January 2025). Installation and loading of the package can be performed using:
@@ 2.4 Detection of Concurrent Mental Health & Substance Use (MHSU) Status – Part (I)
As the third working example, we build on the earlier two to detect the Concurrent Mental Health and Substance Use (MHSU) status for clients diagnosed with psychotic, mood, anxiety, or neurocognitive disorders in combination with alcohol, fentanyl, cannabis, or cocaine consumption-related disorders. We require a minimum of one hospital visit and one physician visit within two months (60 days) for each of MH and SU; the maximum time span between MH and SU statuses is set to one year (365 days). Note that our total dataset is 363 days, shorter than 365 days. The code is below; results appear in Figure 3c:
@@ 3 A Simulation Study
This section introduces a simulation study, encompassing its data and analysis, and is organized into four key parts. Section 3.1 outlines the characteristics of the simulated real-world dataset. Section 3.2 examines how varying the maximum time span within MH status and within SU status affects the frequency of MH, SU, and MHSU diagnoses. Section 3.3 investigates the influence of the number of required hospital and physician visits on MH, SU, and MHSU detection, assuming a fixed maximum time span for MH and SU. Finally, Section 3.4 explores the effect of varying the maximum time span between MH and SU on the number of MHSU diagnoses, assuming fixed maximum time spans within MH and SU.
@@ 3.2 The Impact of Maximum Time Span Within Mental Health and Within Substance Use
Figure 4 presents the results. As $x$ increases, the number of patients detected rises for MH, SU, and MHSU, eventually reaching an asymptote at 125 for MH, 115 for SU, and 90 for MHSU.
@@ 3.3 The Impact of Number of Hospital Visits and Medical Service Physician Visits
Figure 5: Frequency of patients’ diagnosis status vs. number of hospital visits, with a hospital:physician ratio = 1:2 and t_MHSU=365.
@@ 3.4 The Impact of Maximum Time Span for Concurrent Diagnosis
Figure 6: Frequency of concurrent MHSU diagnosis vs. maximum time span between MH and SU (t_MHSU), for t_MH = t_SU = {14,21,28} days.
@@ 4.2 Limitations and Future Work
This work has several limitations, each of which presents opportunities for future extensions. First, the package currently generates outputs as dataframes only, without providing summary frequency statistics for mental health, sub-stance use, and their concurrent statuses. Calculating these statistics requires additional programming. Second, the window lag in the function MHSU_status_broad() is fixed at one day (as illustrated in Figure 2). Adding a parameter to specify the length of this window lag would enhance the flexibility of the detection process for researchers. Finally, for large-scale databases, it may be more efficient to partition the input dataset into multiple disjoint subsets and apply the MHSU_status_broad() function to each subset separately to manage the extensive outputs. Introducing an additional parameter within the function to automate the partitioning of the input data ($inputdata$) would further optimize the computational process.
@@ 5 Conclusions
Supplementary Material. The following supporting information can be downloaded at: www.arxiv.com/xxx/s1, CMHSU R package materials.

### v2

@@ 1.2 North American Healthcare Administrative Databases
Country Level Hospital Data Physician Data Canada Federal Discharge Abstract Database (DAD) National Ambulatory Care Reporting System (NACRS) CIHI Physician Claims Data Canadian MIS Database (CMDB) Province MED-ÉCHO (Quebec) data Ontario Health data Alberta Health data Medical Services Plan (MSP) Data (BC) Alberta Health Physician Claims Régie de l’assurance maladie du Québec (RAMQ) Ontario Health Insurance Plan (OHIP) Territory Yukon Hospital Corporation Dis-charge Data Northwest Territories Health Authority Records Yukon Health Care Insurance Plan Claims Northwest Territories Medical Billing Data United States Federal National Inpatient Sample (NIS) Medicare Provider Analysis and Review (MEDPAR) Veterans Affairs (VA) Inpatient Data Medicare Physician/Supplier Part B Claims Data National Ambulatory Medical Care Survey (NAMCS) Veterans Affairs (VA) Outpatient Data State State Inpatient Databases (SID) California Office of Statewide Health Planning and Development (OSHPD) New York Statewide Planning and Research Cooperative System (SPARCS) State Medicaid Claims Data All-Payer Claims Databases (APCDs: e.g., Massachusetts, Colorado) Territory Guam Memorial Hospital Authority Records American Samoa Dept. of Health Inpatient Data Guam Medicaid and CHIP Claims Data American Samoa Medicaid Claims1
@@ 2.1 The Background and Installation
(accessed on 10 January 2025). Installation and loading of the package can be performed using R commands:
@@ 2.4 Detection of Concurrent Mental Health & Substance Use (MHSU) Status – Part (I)
As the third working example, we build on the earlier two to detect the Concurrent Mental Health and Substance Use (MHSU) status for clients diagnosed with psychotic, mood, anxiety, or neurocognitive disorders in combination with alcohol, fentanyl, cannabis, or cocaine consumption-related disorders. We require a minimum of one hospital visit and one physician visit within two months (60 days) for each of MH and SU; the maximum time span between MH and SU statuses is set to one year (365 days). It is important to note that the total dataset time span is 363 days, which is slightly shorter than the specified maximum span of 365 days. The R code is below; results appear in Figure 3c:
@@ 3 A Simulation Study
This section introduces a simulation study, encompassing its data and analysis, and is organized into four key parts. Section 3.1 outlines the characteristics of the simulated real-world dataset. Section 3.2 examines how varying the maximum time span within MH status and within SU status affects the frequency of MH, SU, and MHSU diagnoses, given the assumption of a constant number of hospital and medical service physician visits. Section 3.3 investigates the influence of the number of required hospital and physician visits on MH, SU, and MHSU detection, assuming a fixed maximum time span for MH and SU. Finally, Section 3.4 explores the effect of varying the maximum time span within MH, within SU, and between MH and SU on the number of MHSU diagnoses, assuming fixed number of hospital and medical service physician visits.
@@ 3.2 The Impact of Maximum Time Span Within Mental Health and Within Substance Use
Figure 4 presents the results. As the maximum required time interval for diagnosis increases, the number of patients detected rises for MH, SU, and MHSU, eventually reaching an asymptote at 125 patients for MH, 115 patients for SU, and 90 patients for MHSU.
@@ 3.3 The Impact of Number of Hospital Visits and Medical Service Physician Visits
Figure 5: Frequency of patients’ diagnosis status vs. number of hospital visits, with a hospital:physician count ratio = 1:2 and t_MHSU=365.
@@ 3.4 The Impact of Maximum Time Span for Concurrent Diagnosis
Figure 6: Frequency of concurrent MHSU diagnosis vs. maximum time span between MH and SU (Time=t_MHSU), for t_MH = t_SU = {14,21,28} days.
@@ 4.2 Limitations and Future Work
This work has several limitations, each of which presents opportunities for future extensions. First, the package currently generates outputs as dataframes only, without providing summary frequency statistics for mental health, sub-stance use, and their concurrent statuses. Calculating these statistics requires additional R programming. Second, the window lag in the function MHSU_status_broad() is fixed at one day (as illustrated in Figure 2). Adding a parameter to specify the length of this window lag would enhance the flexibility of the detection process for researchers. Finally, for large-scale databases, it may be more efficient to partition the input dataset into multiple disjoint subsets and apply the MHSU_status_broad() function to each subset separately to manage the extensive outputs. Introducing an additional parameter within the function to automate the partitioning of the input data ($inputdata$) would further optimize the computational process.
@@ 5 Conclusions
Supplementary Material. The following supporting information can be downloaded at: https://cran.r-project.org/package=CMHSU, CMHSU R package materials.